# O90 O1em m #3

### v1

@@ 3.3 Mapping to Quantum Competencies
The three tables 1, 2 and 3 show which course teaches which competence up to which level, with the domains of possible knowledge and skills taken from the European Competence Framework for Quantum Technologies [1].
@@
Table 1: Competence category: Concepts and Foundation
@@
Table 2: Competence category: Quantum communication and networks
@@
Conventional and PQC, combined cryptographic approaches
A1 A1
Quantum teleportation, Bell state measurement
A1 A1
Security proof, side-channel-attacks
A1
@@
Secure keys, e.g. for QKD
A1 A1 A1
Random numbers for algorithms, e.g. online gambling
A1
@@
QKD basic protocols, e.g. BB84, B92, E91
A1 A2
Measurement-device independent QKD and device independent QKD
A1
Quantum key management systems, QKD modules
A1
@@
Secure access to cloud-based quantum computing, delegated quantum computing
A1
@@
Quantum network nodes, memories and switches
A1
Quantum repeaters, entanglement swapping, entanglement purification
A1 A1
Free-space communication
A1
Quantum internet applications)
Full quantum communication network, QKD trusted node networks (secure data transfer)
A1
Table 3: Competence category: Quantum Computing and Simulation
@@
Qubits, quantum gates, …
A1 A2 A1 A2
Circuit design, notation, matrix representation
A1 A1 A1 A2
Basic quantum programming techniques
A1 A1 A2
@@
Quantum assembler languages and software development kits, quantum circuit simulation
A1 A2
Quantum compilers,high-level programming with pre-defined subroutines
A2
Cloud platforms
A1 A2
Quantum error correction
A1
@@
Quantum amplitude amplification
A2
QFT, hidden subgroup finding
A1
Quantum phase estimation
A1 A2
Quantum linear algebra subroutines, quantum singular value decomposition
A1
@@
Numer theory and factorisation
A1
Oracular algorithms and database search
A1 A1 A2
Linear algebra (e.g. HHL-algorithm)
A1
Quantum optimisation
A1
Quantum machine learning, quantum neural networks
A1
Quantum simulation algorithms
A1
NISQ-Algorithms, VQE, QAOA
A1
@@
Data security and cryptography
A1 A1
@@ 4 Evaluation of the data collected
Across all courses in the curriculum under consideration, there are till January 2024 a total of 17157 course attendances of 7413 distinct participants. Among those 6232 participants were already registered on OpenHPI; 1181 created a new account on the same day they enrolled in their first course.
4.1 RQ 1: Did the MOOCs address a broader audience with different professional and personal characteristics?
In this section, the personal and professional background of the participants is analyzed. The information provided voluntarily by the participants is used for this purpose. Thus, the information relates to about 5700 of 17147 registrations, and participants who attended several courses may have answered more than once. Nevertheless, the analyses provide good indications of the spectrum of participants present and of frequently occurring characteristics.
@@
It turns out that the typical participant in a course is between 50 and 59 years old, male, employed in IT, and has a diploma or master’s degree. These characteristics are mentioned in between 30% and 80% of registrations. But often there is a wide variety of answers for the non-typical participants. The courses were therefore also attended by hundreds to thousands of participants who are female, do not work in IT, or do not have a university degree.
Figures (2) - (6) provide detailed information: In Figure (2) we show the distribution of age groups over all participants, without double countings of participants. Unfortunately, the majority (4747 participants) did not disclose their age. Of those who revealed their age, a majority of participants are in the 50-59 age group.
@@
Figure (4) shows the distribution of genders. At over 80%, the proportion of men is even higher than is typically observed in STEM subjects, where the proportion of men is regularly around 2/3.
Figures (4) - (6) reveal information about professional characteristics. Over 60% of survey participants described their employment status as employed, followed by retired people (approx. 17%) and students at universities and others (approx. 7% each). Unsurprisingly, around 58% of participants categorize themselves in the IT sector, followed by other (around 16%), education (around 14%), and engineering (around 12%). However, it was also possible to reach people from administration and marketing/sales. In terms of professional qualifications, the MA degree dominates with approx. 54%, followed by a Bachelor’s degree (approx. 18%) and a doctorate (approx. 17%). Around 15% of respondents stated that they had a non-academic professional qualification.
@@
4.2 RQ 2: Did the participants follow given learning paths and other learning recommendations?
@@
When analyzing the items attended per course, after deducting the no-shows who were only enrolled and did not attend any items at all, the largest group of participants attended 75% to 100% of the items. The second largest group only attended less than 25% of the items. It turns out that only a small percentage of participants visited an average number of items between 26% and 75%.
@@
Figures (7) - (10) show the details. Figure (7) shows that many participants specifically selected one or a few courses and did not decide to complete the entire program with all courses. The number of participants attending one course, two courses, .., nine courses can be seen.
@@
Figure (8) shows the registrations per course. Summed up over all nine courses, there are 17157 registrations until January 2024. 12468 are course learners and 4689 are self-learners. The courses continue to be open for self-learners, so the number of self-learners is still increasing.
Figure 8: Registrations per course, subdivided to course learners and self-lerners.
Figure (9) displays for each course initially how many participants visited how many items. Participants are divided into 5 groups: the no-shows (participants who registered but did not visit any course item), and then the 4 groups of participants who visited between 1% and 25%, between 26% and 50%, between 51% and 75%, and more than 75% of the items. In all courses, the middle participant groups, those with shares between 26% and 75%, were by far the smallest groups. The other groups are roughly equal in size, with variations between individual courses.
@@
Figure (10 ) shows which participants took which learning paths. Here, only course participations with at least 25% of visited items are considered, these are 2036 participants. The most visited learning path consists of the course Intro 1 only, with 720 participants. This is followed by the learning path Intro 1 and Crypto 1 with 316 participants, and then the learning path with all courses with 119 participants. Only in exceptional cases course selections outside of the recommended learning paths were made.
@@
4.3 RQ 3: Were the desired competencies acquired by the course participants, and did using the recommended learning items contribute to the learning success?
To check the success of course participation, we split the 17157 data records into 12468 data records from course learners (CLs) and 4689 data records from self-learners (SLs). The distinction between course learners and self-learners was made in the data records based on the enrolment date for the course.
@@
The learning success of course learners is derived from the result of the final examination. This is passed if 50% of the questions are answered correctly.
It can be seen that of the 12468 course learners, 4550 are no-shows, 3145 did not reach the 50% items threshold and 4773 successfully completed the course. This corresponds to 38% of all course learners.
Figure (11) shows the correlation between the learning success (proportion of the total score achieved including the final exam) and the proportion of the learning items attended. While only a few participants achieved a high overall score despite having attended fewer learning items, the majority showed a clear positive correlation between the two variables. A simple linear regression model of the score achieved with the explanatory variable proportion of items visited achieves an $R^{2}=0.27$. This underlines the significance of the learning materials visited for course success.
Figure 11: Points percentage (=total score) of the course learners vs. percentage of items visited.
In Fig(13) we show that the mean of the number of items visited is significantly larger for the top performers of the courses (top 5%, top 10% and top 20% of the participants) compared with the remaining participants. Here ”topx” means that a participant has achieved one of the x highest percentage results (i.e. the result of a top 5 performer is not necessarily better than that of 95% of the field of participants). The top performers visited more than 92% of the items on average.
@@
Figure 13: Percentage of the items by Top Performers (Top 5, Top10, Top 20) and the other course-learners.
Fig(14) gives an overview and the details on course registrations for course learners and self-learners, the no-shows, the unsuccessful, and the successful participants. Across all courses and participants, 5912 completed the course; in 11245 cases the courses were not completed. It should be noted here that individual participants are counted separately in each course they attended.
Figure 14: Divided between course and self-learners the number of no-shows, unsuccessful, and successful participants are displayed.
@@ 5 Conclusions and topics for further research
RQ 1:
Did the MOOCs address a broader audience with different professional and personal characteristics?
RQ 2:
Did the participants follow the given learning paths and other learning recommendations?
RQ 3:
Were the desired competencies acquired by the participants in the courses, and did using the recommended learning items contribute to the learning success?
@@
RQ1: Our analyses show that a broad audience attended the quantum MOOCs. All age groups were represented, with the majority being over forty. Most participants had an IT background, but participants from non-technical professions are also a considerable part of the audience. The majority of participants has a university degree. As is often the case in technical and scientific fields, male participants also predominated in our courses. The analyses show that people not directly associated with a university are also interested in quantum computing. A broader audience with different professional and personal characteristics could therefore be addressed.
@@
RQ2: As the field of quantum computing is very broad, various learning paths were suggested to the participants. With the MOOC platform’s help, the participants’ learning behavior could be tracked in detail. The analyses show that the recommended learning paths were often followed. Most participants attended an introductory course in each subject area (Intro 1, Crypto 1, Algo 1). Depending on their interests, they remained loyal to the topic and completed more advanced courses. A not inconsiderable number of participants also attended courses on several topics. More than 140 people completed all nine core MOOCs. The data shows that the recommenadtion of learning paths was highly accepted. Thus, there seems to be evidence that it is imporatant to suggest learning paths to participants for better orientation when dealing with complex topics.
@@
RQ3: The generally recognized European Reference Framework for Quantum Technologies was used to select material. All content taught and tested in the examinations could be assigned to categories in the reference framework. Over 80% of exam participants passed the final exam, 50% even with a good to excellent result. It also shows that the success rate correlates with the number of learning materials consumed (videos, self-study tests, forum discussions). It can be stated that the learning elements provided by the MOOC platform are sufficient to teach participants the necessary skills.
In conclusion, one can say that the quantum MOOCs at OpenHPI were a big success and up to January 2024 reached more than 7400 participants from different age and professional groups. The MOOCs help to bring the topic of quantum computing closer to a broad audience outside of universities and research departments. Therefore, the topic has reached a broader audience.
@@
In addition to the results discussed in the present paper, further research questions can be investigated based on the available data. It would be interesting to see, how exactly the participants worked with the video tutorials. For example, were the videos paused in between to better understand complicated details? Did the sometimes detailed discussions in the forum motivate the participants to engage with the topic? Another interesting question could deal with possible specific learning differences between male and female participants. More generally, the learning behavior of different subgroups could be analyzed in relation to age, professional situation, or educational attainment. Since the amount of data is large, the subgroups usually consist of hundreds of participants with respect to various characteristics, so that statistical analyses are meaningful.
@@
Acknowledgements We would like to thank the Hasso-Plattner-Institut für Digital Engineering gGmbH and in particular Christoph Meinel and Martin van Elten for providing the data on which this study is based.

### v2

@@ 3.3 Mapping to Quantum Competencies
The three Tables 1, 2 and 3 show which course teaches which competence up to which level, with the domains of possible knowledge and skills taken from the European Competence Framework for Quantum Technologies [1]. The framework defines six proficiency levels: A1 Awareness, A2 Literacy, B1 Utilisation, B2 Investigation, C1 Specialisation, and C2 Innovation.
@@
Table 1: Competence Domains: Concepts and Foundation
@@
Table 2: Competence Domains: Quantum communication and networks
@@
Conventional and PQC, combined cryptographic approaches A1 A1
Quantum teleportation, Bell state measurement A1 A1
Security proof, side-channel-attacks A1
@@
Secure keys, e.g. for QKD A1 A1 A1
Random numbers for algorithms, e.g. online gambling A1
@@
QKD basic protocols, e.g. BB84, B92, E91 A1 A2
Measurement-device independent QKD and device independent QKD A1
Quantum key management systems, QKD modules (full devices) A1
@@
Secure access to cloud-based quantum computing, delegated quantum computing A1
@@
Quantum network nodes, memories and switches A1
Quantum repeaters, entanglement swapping, entanglement purification A1 A1
Free-space communication A1
Quantum internet applications
Full quantum communication network, QKD trusted node networks (secure data transfer) A1
Table 3: Competence Domains: Quantum Computing and Simulation
@@
Qubits, quantum gates, … A1 A2 A1 A2
Circuit design, notation, matrix representation A1 A1 A1 A2
Basic quantum programming techniques A1 A1 A2
@@
Quantum assembler languages and software development kits, quantum circuit simulation A1 A2
Quantum compilers,high-level programming with pre-defined subroutines A2
Cloud platforms A1 A2
Quantum error correction A1
@@
Quantum amplitude amplification A2
QFT, hidden subgroup finding A1
Quantum phase estimation A1 A2
Quantum linear algebra subroutines, quantum singular value decomposition A1
@@
Numer theory and factorisation A1
Oracular algorithms and database search A1 A1 A2
Linear algebra (e.g. HHL-algorithm) A1
Quantum optimisation A1
Quantum machine learning, quantum neural networks A1
Quantum simulation algorithms A1
NISQ-Algorithms, VQE, QAOA A1
@@
Data security and cryptography A1 A1
@@ 4 Evaluation of the data collected
Across all courses in the curriculum under consideration, there are till January 2024 a total of 17,157 course attendances of 7413 distinct participants. Among those 6232 participants were already registered on OpenHPI; 1181 created a new account on the same day they enrolled in their first course.
4.1 RQ 1: What kind of audience attended the MOOCs? What are the specific backgrounds and interests of the participants?
In this section, the personal and professional background of the participants is analyzed. The information provided voluntarily by the participants is used for this purpose. Thus, the information relates to about 5700 of 17,147 registrations, and participants who attended several courses may have answered more than once. Nevertheless, the analyses provide good indications of the spectrum of participants present and of frequently occurring characteristics.
@@
It turns out that the typical participant in a course is between 50 and 59 years old, male, employed in IT, and has a university degree. These characteristics are mentioned in between 30% and 80% of registrations. But often there is a wide variety of answers for the non-typical participants. The courses were therefore also attended by hundreds to thousands of participants who are female, do not work in IT, or do not have a university degree.
Figures 2-6 provide detailed information: In Fig. 2 we show the distribution of age groups over all participants, without double countings of participants. Unfortunately, the majority (4747 participants) did not disclose their age. Of those who revealed their age, a majority of participants are in the 50-59 age group.
@@
Figure 4 shows the distribution of genders. At over 80%, the proportion of men is even higher than is typically observed in STEM subjects, where the proportion of men is regularly around 2/3.
Figures 4-6 reveal information about professional characteristics. Over 60% of survey participants described their employment status as employed, followed by retired people (approx. 17%) and students at universities and others (approx. 7% each). Unsurprisingly, around 58% of participants categorize themselves in the IT sector, followed by other (around 16%), education (around 14%), and engineering (around 12%). However, it was also possible to reach people from administration and marketing/sales. In terms of professional qualifications, the MA degree (this also includes the former German diploma degrees) dominates with approx. 54%, followed by a Bachelor’s degree (approx. 18%) and a doctorate (approx. 17%). Around 15% of respondents stated that they had a non-academic professional qualification.
4.2 RQ 2: What specific topics or subject areas did the participants choose when selecting the courses offered? Which topics are the participants particularly interested in?
@@
When analyzing the items attended per course, after deducting the no-shows who were only enrolled and did not attend any items at all, the largest group of participants attended 75% to 100% of the items. The second-largest group only attended less than 25% of the items. It turns out that only a small percentage of participants visited an average number of items between 26% and 75%.
@@
Figures 7-10 show the details. Figure 7 shows that many participants specifically selected one or a few courses and did not decide to complete the entire program with all courses. The number of participants enrolling in one course, two courses, …, and all nine courses can be seen.
@@
Figure 8 shows the registrations per course. Summed up over all nine courses, there are 17,157 registrations until January 2024. 12,468 are course learners and 4689 are self-learners. The courses continue to be open for self-learners, so the number of self-learners is still increasing.
Figure 8: Registrations per course, subdivided to course learners and self-learners.
Figure 9 displays for each course initially how many participants visited how many items. Participants are divided into 5 groups: the no-shows (participants who registered but did not visit any course item), and then the 4 groups of participants who visited between 1% and 25%, between 26% and 50%, between 51% and 75%, and more than 75% of the items. In all courses, the middle participant groups, those with shares between 26% and 75%, were by far the smallest groups. The other groups are roughly equal in size, with variations between individual courses.
@@
Figure 10 shows which participants took which learning paths. Here, only course participations with at least 25% of visited items are considered, these are 2036 participants. The most visited learning path consists of the course Intro 1 only, with 720 participants. This is followed by the learning path Intro 1 and Crypto 1 with 316 participants, and then the learning path with all courses with 119 participants. Only in exceptional cases course selections outside the recommended learning paths were made.
@@
4.3 RQ 3: Which competencies could be acquired by the participants through the course concept? To what extent did the course concept support the participants?
To check the success of course participation, we split the 17,157 data records into 12,468 data records from course learners (CLs) and 4689 data records from self-learners (SLs). The distinction between course learners and self-learners was made in the data records based on the enrolment date for the course.
@@
The learning success of course learners is derived from the result of the final examination. This is passed if 50% of the questions are answered correctly. These participants get a “Record of Achievement”. All participants who visited at least 50% of the items in a course get a “Confirmation of Participation”.
It can be seen that of the 12,468 course learners, 4550 are no-shows, 3145 did not reach the 50% items threshold and 4773 successfully completed the course. This corresponds to 38% of all course learners.
To gain more insights on which the points percentage (total score) of the course learners may depend, we model this with a simple regression model where the points percentage depends linearly on the percentage of items visited, see Fig. 11. Here, besides the data points, the linear model with slope and intercept as well as the $R^{2}$ can be seen. While only a few participants achieved a high overall score despite having attended fewer learning items, the majority showed a clear positive correlation between the two variables as seen by the value of $R^{2}=0.27$. Note, that as a global model, the linear regression interpolates smoothly between different regimes of the data.
Aggregated over all courses, 3299 participants tried the final exams and are included in this analysis. This underlines the significance of the learning materials visited for course success.
Figure 11: Points percentage (=total score) of the course learners vs. percentage of items visited. 3299 data points are included.
When dividing the dataset into the three course blocks, Intro 1 to 3, Crypto 1 to 3, and Algorithms 1 to 2/3, we observe a qualitatively similar dependence on points percentage and items visited percentage: Participants using the content of the courses more actively performed better in the final exam. This is shown in Figs. 12 to 14.
Figure 12: Courses Intro 1, Intro 2 and Intro 3: Points percentage (=total score) of the course learners vs. percentage of items visited. 1667 data points are included.
Figure 13: Courses Crypto 1, Crypto 2 and Crypto 3: Points percentage (=total score) of the course learners vs. percentage of items visited. 1051 data points are included.
Figure 14: Courses Algo 1 and Algo 2/3: Points percentage (=total score) of the course learners vs. percentage of items visited. 581 data points are included.
For the course learners, we can furthermore relate the success of the final exam to the items visited. Figure 15 gives an overview on the results achieved in the final exams. In Fig. 16 we show that the mean of the number of items visited is significantly larger for the top performers of the courses (top 5%, top 10% and top 20% of the participants) compared with the remaining participants. Here “topx” means that a participant has achieved one of the x highest percentage results. Surprisingly, the Top 5 performers attended slightly fewer items than the Top 10 and Top 20 performers. This may be due to the fact that some quantum professionals also attended the courses to relate them to their professional knowledge or to use the structure of the courses. Overall, the top performers visited more than 92% of the items on average.
@@
Figure 16: Percentage of the items visited by Top Performers (Top 20, Top 10, Top 5) and the other course-learners.
Figure 17 gives an overview and the details on course registrations for course learners and self-learners, the no-shows, the unsuccessful, and the successful participants. Across all courses and participants, 5912 completed the course; in 11,245 cases the courses were not completed. It should be noted here that individual participants are counted separately in each course they attended.
Figure 17: Divided between course and self-learners the number of no-shows, successful and unsuccessful participants are displayed.
@@ 5 Conclusions and topics for further research
  1. [RQ 3:]
  2. RQ 1:
    What kind of audience attended the MOOCs? What are the specific backgrounds and interests of the participants?
  3. RQ 2:
    What specific topics or subject areas did the participants choose when selecting the courses offered? Which topics are the participants particularly interested in?
  4. RQ 3:
    Which competencies could be acquired by the participants through the course concept? To what extent did the course concept support the participants?
@@ 4. RQ 3:
RQ1: Our analyses show that the quantum MOOCs were attended by an audience with different backgrounds and interests. All age groups were represented, with the majority being over forty. Most participants had an IT background, but participants from non-technical professions are also a considerable part of the audience. The majority of participants have a university degree. As is often the case in technical and scientific fields, male participants also predominated in our courses. The analyses show that people not directly associated with a university are also interested in quantum computing. An enlarged audience with different professional and personal characteristics could therefore be addressed.
@@
RQ2: As the field of quantum computing is very broad, various learning paths were suggested to the participants. With the MOOC platform’s help, the participants’ learning behavior could be tracked in detail. The analyses show that the recommended learning paths were often followed. Most participants attended an introductory course in each subject area (Intro 1, Crypto 1, Algo 1). Depending on their interests, they remained loyal to the topic and completed more advanced courses. A not inconsiderable number of participants also attended courses on several topics. More than 140 people completed all nine core MOOCs. The data shows that the recommendation of learning paths was highly accepted. Thus, there seems to be evidence that it is important to suggest learning paths to participants for better orientation when dealing with complex topics.
@@
RQ3: The generally recognized European Competence Framework for Quantum Technologies [1] was used to select material. All content taught and tested in the examinations could be assigned to categories in the reference framework. Over 80% of exam participants passed the final exam, 50% even with a good to excellent result. It also shows that the success rate correlates with the number of learning materials consumed (videos, self-study tests, forum discussions). It can be stated that the learning elements provided by the MOOC platform are sufficient to teach participants the necessary skills.
In conclusion, one can say that the quantum MOOCs at OpenHPI were a big success and up to January 2024 reached more than 7400 participants from different age and professional groups. The MOOCs help to bring the topic of quantum computing closer to an audience outside of universities and research departments. The MOOCs are a good opportunity for anyone interested in quantum computing to delve deeper into the topics. The structure and content of the courses might also be a good point of reference and starting point for persons who would like to hold quantum computing courses. Depending on the desired learning objectives, teachers can build their own courses along the proposed learning paths, since the learning paths provide a self-contained arrangement of quantum computing topics. The teachers then can deepen quantum topics according to their specialization, and insert additional topics according to their preferences.
As the courses are in high demand, the openHPI has now started to provide the videos with English subtitles to make them accessible to a non-German-speaking audience. However, as the slides shown are still in German, the reach of the courses is likely to remain limited. In hindsight, it would have been better to formulate the slide content directly in English. The project schedule (start, milestones, end) required by the project organizer meant that the timing of the courses was not always optimal. Some courses were held very close together, while other courses were sometimes several months apart. When organizing an extensive series of courses, care should be taken to ensure that the courses are evenly distributed, giving the participants enough time to catch up on their work, but that the breaks between the courses are not too long.
The experience gained from running the MOOCs shows that many concepts of quantum computing can be explained to an interested audience without complex mathematical and physical descriptions. Presenting the various topics using examples and applications that are closer to everyday life has proved useful. Mathematical descriptions were only used when there was an intuitive understanding based on the application. Such an approach lowers the entry barrier to the topic and leads to less frustration among learners. Mathematical and physical descriptions and concepts should only be introduced and used once a basic understanding and curiosity have been aroused.
@@
In addition to the results discussed in the present paper, further research questions can be investigated based on the available data. It would be interesting to see, how exactly the participants worked with the video tutorials. For example, were the videos paused in between to better understand complicated details? Did the sometimes detailed discussions in the forum motivate the participants to engage with the topic? Another interesting question could deal with possible specific learning differences between male and female participants. More generally, the learning behavior of different subgroups could be analyzed in relation to age, professional situation, or educational attainment. The professional and personal backgrounds of the Top Performers could also be analyzed in more detail. Since the amount of data is large, the subgroups usually consist of hundreds of participants with respect to various characteristics, so that statistical analyses are meaningful.
@@
Acknowledgements The authors would like to thank the Hasso-Plattner-Institut für Digital Engineering gGmbH and in particular Christoph Meinel and Martin van Elten for providing the data on which this study is based, and the unknown reviewers for their valuable hints.
@@
Appendix A The “Quantum Cube” for visualizing entanglement in multi-qubit-registers and the effect of quantum gates
In the introductory MOOCs of the curriculum, the “quantum cube” is used throughout to illustrate quantum register states and the effect of quantum gates.
The model was developed by Just [29, 30] and since then has been studied and used as an intuitive tool for visualizing quantum entanglement and quantum computation [31, 32, 33]. The quantum cube represents the state of an n-qubit register by providing one dimension in n-dimensional Euclidean space to each qubit. It then utilizes the standard n-dimensional cube with corners $(0,\ldots,0)$ up to $(1,\ldots,1)$. For a given register state, it places the amplitudes of the basis states (in any suitable representation of numbers, i.e. digits, arrows, squares, cubes, or circles) at the corresponding corners of the n-dimensional cube. See Fig. 18 for an example with two qubits, and the effect of Pauli-X-gate operating on the first of them.
The application of quantum gates to a quantum register can generally be visualized as a swapping, rotation, or mixing of the numbers at the corners of the cube. This has been done by Just [29, 30] for the teleportation algorithm, see Fig. 19.
For the purpose of the quantum MOOCs at openHPI, the cube model for the first time was applied to more advanced quantum algorithmic features, such as quantum error correction, see Fig. 20, quantum oracles, see Fig. 21, and quantum phase kickback, see Fig. 22. Based on the MOOCs, the visualization of some abstract algorithmic concepts with the quantum cube model is summarized and demonstrated in [44].
Figure 18: Visualizing the state $\sqrt{\frac{1}{16}}e^{i290^{o}}|00\rangle+\sqrt{\frac{4}{16}}i|01\rangle+\sqrt{\frac{5}{16}}|10\rangle+\sqrt{\frac{6}{16}}e^{i200^{o}}|11\rangle$ of a 2-qubit-register and the effect of Pauli-X operating on the first qubit with the quantum cube model. The first qubit is provided with the left-right direction in 2-dimensional space, the second qubit with the down-up direction. The amplitudes are placed at the corresponding corners of the n-dimensional unit cube, which in 2 dimensions is the unit square. The representation of the complex amplitudes here follows Feynman’s QED [45]. Pauli-X on the first qubit interchanges the amplitudes along the left-right lines of the “cube”. The resulting state is $\sqrt{\frac{5}{16}}|00\rangle+\sqrt{\frac{6}{16}}e^{i200^{o}}|01\rangle+\sqrt{\frac{1}{16}}e^{i290^{o}}|10\rangle+\sqrt{\frac{4}{16}}i|11\rangle$.
Figure 19: The CNOT-Hadamard-Measurement-steps of the teleportation algorithm as in [29]. For didactical reasons, only real amplitudes are considered. Green squares indicate positive amplitudes, red squares indicate negative amplitudes. The state of a register with three qubits is visualized in a 3-dimensional cube. Each qubit is provided with one dimension in space: The first qubit with the left-right dimension, the second qubit with the down-up direction, the third qubit with the front-back direction.
Figure 20: The decoding in bit flip correction code, visualised with the quantum cube model. The CNOT gate with the first qubit as control-bit is applied two times, first with the second qubit as target-bit (down-up interchangement of amplitudes when first qubit equals $|1\rangle$), then with the third qubit as target-bit (front-back interchangement of amplitudes when first qubit equals $|1\rangle$). It can be seen how the amplitudes $\alpha$ and $\beta$ of the first qubit (left-right direction) are rearranged to one specific state of the syndrome qubits (down-up and front-back direction).
Figure 21: A quantum oracle for a boolean function of three bits, in this example $f(x_{1},x_{2},x_{3})=1$ if and only if $(x_{1},x_{2},x_{3})$ is $(0,0,0)$ or $(1,0,1)$ or $(1,1,1)$. The oracle $U_{f}$ in general interchanges the amplitudes of the $|x_{0}\rangle$ and $|x_{1}\rangle$ states, if and only if $f(x)=1$. The initial state in this figure is prepared as it is done in the algorithms of Deutsch-Jozsa and Bernstein-Vazirani
Figure 22: Illustrating quantum phase kickback für a controlled-U. The unitary U operating on two qubits changes the amplitudes of a register from orange to grey (which could mean anything), but in the case of an eigenstate, this means turning all of them by the same angle. Controlled-U operating on three qubits applies U to the states where the control-qubit is 1, hence to those at the left-sided square of the cube. Applied to an eigenstate, controlled-U just turns the amplitudes of the left-sided square of the cube all by the same angle.